I need help solving the following questions, I cannot do them for the life of me.
14 x - 4(9 x - 40) = 345 (solving for x)
5.25 x^4 – 8 = 15 (solve for x)

Answer by mananth(16946)

14 x - 4(9 x - 40) = 345 (solving for x)
14x-36x+160=345
-22x+160=345
-160
-22x=345-160
-22x=185
/-22
x=185/-22
5.25 x^4 – 8 = 15 (solve for x)

25x^4-8=15
25x^4=23
/25
x^4= 23/25
x^2= +/- sqrt(23)/5
x= +/- 4th root of 23/sqrt(5)